# UCINet0: A Machine Learning based Receiver for 5G NR PUCCH Format 0

### v1

@@ III-B2 Real-Time Over-The-Air Datasets
We use hardware captures derived from the state-of-the-art 5G testbed at IIT Madras [7] for more realistic testing. The hardware captures help represent channel states and hardware impairments not included in the training dataset. The setup (see Figure 3) consists of an N5182B Vector Signal Generator (VSG) for transmitting the 5G signal at a center frequency of $3.49986$ GHz (one of the sub-6 GHz channel raster in the n78 band). In our setup, the VSG acts like a UE, and it uses a commercial omnidirectional wideband monopole antenna to transmit the signals. The VSG is connected to the antenna through 2 SMA cables with $1.9$ dB wire loss each. Here, commercial UEs could not be used because it is not easy to extract the actual transmitted information, i.e., ground truth labels for the UCI transmissions made by the UE. Also, the dynamic nature of the allocations made by the L2 scheduler makes it difficult to capture the data at the gNB for that particular time duration.
@@
A multi-channel 5G Remote Radio Head (RRH) with a dual-polarized antenna at the gNB receives the signals from over the air (for the purpose of the paper, we utilize only one antenna and one transceiver chain). The RRH operates in the n78 band with $100$ MHz bandwidth. It is ORAN-compliant and follows the 7.2b split as defined in [17]. Other receiver components of the RRH include an in-house Low Noise Amplifier (LNA) with 60dB gain at the receiver front end and an ADRV 9009 RF transceiver. We place the transmitter antenna one meter away from the receiver. The PUCCH Format 0 signal is transmitted from the VSG through an antenna, over the air, and then received at the RRH antenna, followed by the LNA and the transceiver. The signal out of the 16-bit ADC is then collected and used for testing.
@@
Figure 3: IIT-Madras 5G testbed setup with the Remote Radio Head used as a receiver (gNB) and the VSG used as a transmitter (Emulated UE).
Figure 4: UCINet0 Architecture for PUCCH Format 0 decoding with 25 neurons in the input layer, 512 neurons in each of the three hidden layers and 12 neurons in the output layer.
III-C Neural Network Architecture
Given a problem statement, there is no formula for knowing apriori, the best neural network architecture. For most problems, initial architectures are determined based on prior experience and domain knowledge. These architectures are then tuned by experiment. Figure 4 shows the UCINet0 architecture that showed the best performance in our experiments. In subsequent sections, we also comment on the trade-off between high complexity and performance.
III-C1 Neural Network Structure
The input to the NN has two components. The first is the received PUCCH Format 0 signal, and the second is metadata indicating the number of multiplexed UEs that the receiver expects. Since the PUCCH Format 0 signals occupy a single Resource Block, there are 12 received complex samples. We separate the real and imaginary parts of each complex number and concatenate them, resulting in a length 24 real sequence. As shown in Eq. (3), a given PUCCH resource block at the receiver could contain the sum of signals from multiple UEs. To aid the neural network, we also feed the number of multiplexed UEs as a 25th metadata input. The number of multiplexed UEs that a receiver should expect to see is often provided by an L2 scheduler as an upper bound, meaning that the received signal could contain less than or equal to the number of UEs indicated by L2. We denote the difference between the actual number of UEs multiplexed and the number of UEs indicated by L2 as $\Delta$. One example of a scenario where this is observed is when the L2 scheduler allocates multiple UEs to transmit SR, but not all of the UEs have any data to transmit. In such cases, UEs that do not require PUSCH resources do not transmit the SR.
Figure 5: Model accuracy for different FCN architectures at various values of $\Delta$. Here, $\Delta$ represents the maximum metadata offset between the true value and L2’s upper bound for the number of multiplexed UEs. We use these results to arrive at the final UCINet0 architecture presented in this paper.
Owing to the relatively small dimension of the input, our proposed UCINet0 is a Fully Connected Neural Network (FCN). As shown in Figure 4, there are 3 dense layers, each containing 512 neurons. The output layer contains twelve neurons. Each output neuron acts as an indicator for the transmission of the base sequence which is cyclically shifted by $\alpha_{m}$ 3. We arrived at the final UCINet0 architecture of 3 layers and 512 neurons after having evaluated the performance of other architectures with a varying number of layers and neurons. Figure 5 shows the UCI decoding accuracies(defined in IV IV-B) obtained with 20 variations of FCN architectures. We considered combinations of $1$, 22, 33 and $4$ layer models with $64$, $128$, $256$, $512$ and $1024$ neurons.
From Figure 5, we observe the following:
  • Model performance generally improves as we increase the number of neurons from $64$ to $1024$ for any number of layers and for values all of $\Delta$.
  • For $\Delta=0$, both $1024$ and $512$ neurons result in almost the same accuracy values for a given number of layers. For example, $2\text{ layers}\times 1024$ and 2 $2\text{ layers}\times 51$2 result in similar accuracy values.
  • For higher values of $\Delta$, the accuracy drops as the number of layers increases beyond a certain value as the model begins to overfit. For $\Delta=4$ and $1024$ neurons, the accuracy drops beyond 2 layers, whereas for $512$ neurons, the accuracy drops beyond 3 layers.
Hence, considering the balance between the number of layers, number of neurons, and accuracies for various $\Delta$ values, 3 layers with $512$ neurons architecture seems to be the optimal choice.
III-C2 Dropout
Dropout [18] is used during training to ”drop out” or cut off certain neurons and their connections randomly with a certain probability during training. It serves two purposes - Ensemble Learning [19] and Regularization [20]. Ensemble learning helps improve predictive performance by combining the outputs of multiple AI/ML models or multiple instances of the same AI/ML model trained on different subsets of data. However, training a large number of models is extremely prohibitive both in terms of time and computational resources. With dropout, in each backward pass, only a subset of weights are updated, effectively resulting in a new neural network. The subset of weights that are updated varies with each forward and backward pass, thus creating the combined effect of training several neural networks without the significant computational overhead. Dropout also has a regularizing effect on training in that it makes training noisy and, as a result, doesn’t allow neurons to co-adapt. This means that neurons in one layer will learn not to depend on or compensate for other neurons because these very neurons may randomly be dropped out. This encourages neurons to become more independent, thus preventing overfitting. All the models described in this paper use dropout with a probability of 0.5 in each of the hidden layers.
III-C3 Activations and Backpropagation
The hidden layers use a ReLU activation function, and the output layer uses a Sigmoid activation function. The use of the sigmoid activation function is motivated by the fact that in a multi-label classifier, the sum of all the output probabilities need not be 1. Each individual output neuron can be thought to be a part of an independent binary classifier. In other words, since each output neuron corresponds to an $\alpha$ value, the probabilities of each output neuron indicate the confidence with which the NN detects the presence of a certain UE with a certain $\alpha$ value. The output probabilities are used to compute a binary cross entropy loss that is back-propagated using Stochastic Gradient Descent (SGD) with a learning rate of $10^{-2}$ and a momentum parameter of $0.9$.
III-D Training and Testing
We consider received signal data corresponding to 5 SNR values i.e., $0,5,10,15,20$ dB. The training instances are designed such that they encompass all the possible cases of multiplexed UEs including the cases where no UEs are transmitting. During training, we assume that the metadata input always holds the correct number of UEs present in the signal. Furthermore, based on the insights gained from our previous work [6], we train the neural network only using data with a middle SNR of 10dB, rather than all the SNRs. We have found that this model achieves good generalization across the entire range of SNR during inference/testing. 75% of the data is used for training from which a further 30% is used for validation and fine-tuning of hyper-parameters. We train the model for 150 epochs(an epoch is 1 pass of the entire dataset through the neural network). The training and testing was done using Python and TensorFlow. The code is available at [16] for download.
In the inference phase, we test the model using data from all SNRs. We also incorporate the fact that the number of multiplexed UEs in a real-time deployment may not be exactly known. The L2 scheduler in a gNB merely provides a value for the upper bound of the number of UEs a receiver could possibly expect in a given PUCCH allocation. This is usually a consequence of one of the following scenarios:
  • A UE does not transmit any SR if no uplink grants are required because of lack of data into buffer
  • A UE does not transmit feedback (on HARQ resources) if the prior downlink message is not detected.
During testing, we feed this upper bound value as the metadata input to the neural network, which is offset from the true value. To test the robustness of the trained model, we consider various cases of the maximum metadata offset $\Delta$ between the true value and L2’s upper bound for the number of multiplexed UEs. For example, $\Delta=0$ indicates that the number of UEs provided as the metadata input to the NN matches the number of UEs actually transmitting in the given allocation. $\Delta=2$ indicates that for each test instance, the number of UEs provided as the metadata input to the NN could be offset by any value between 0 and 2 with 0 being the best case (no offset) and 2 being the worst. In this paper, we show results for $\Delta=0,2,4$.
Eq. (4) shows the offset added to the true value of the number of multiplexed UEs, $N_{UE}$ to obtain $\tilde{N}_{UE}$ which is the expected number of multiplexed UE at the receiver, as indicated by L2.
$\tilde{N}_{UE}\in[N_{UE},N_{UE}+\Delta]$ (4)
Note that since the maximum number of UEs that can be multiplexed in a given resource block is $12$, the value of $\tilde{N}_{UE}$ is chosen accordingly.
IV Results and Discussion
@@
In this section, we describe the performance results of the UCINet0 Neural Network configuration shown in Figure 4.
IV-A Performace Metrics and their Motivation
@@
We use model test accuracy as the main performance metric. We also show confusion matrices to indicate the distribution of model predictions. Wherever appropriate, we show comparisons of the model performance with the corresponding performance shown by the DFT algorithm.
@@
Accuracy is simply the number of correct predictions in relation to total predictions. Though it is a useful high-level metric to gauge model performance, in cases where the classes in a dataset are unequally distributed or when, as is the case in this paper, there are more than 2 classes, accuracy alone as a metric for an NN classifier does not offer the full picture of model performance. It is helpful to know if the model classifies all classes equally well or if it is more ”confused” by certain classes compared to others. Hence, calculating a confusion matrix gives a better insight into any patterns that may exist in both correct and wrong classifications made by the model. In this paper, we show 3 types of confusion matrices: (1) A multi-label confusion matrix that shows how well each $\alpha$ is predicted; (2) A confusion matrix showing how well the model predicts the correct number of multiplexed UEs, and (3) A confusion matrix showing how each value of $\alpha$ (including the cases when no $\alpha$ is selected) is classified. We present this last metric in the form of a column chart for ease of visualization. These 3 types of confusion matrices provide a sense of interpretability of the inner workings of the NN model.
IV-B Accuracy and Loss
@@
Figure 6(a) shows the training and validation accuracy with respect to the training epochs. Figure 6(b) shows the training and validation loss. Both curves follow the expected pattern for a well-fit model. The accuracy/loss gradually increases/decreases with epoch and eventually converges. We also note that the validation accuracy and loss values are slightly better than the corresponding training values. One possible explanation for this could be the use of dropout. As stated above, a dropout with probability 0.5 drops neurons $50\%$ of the time during training. However, during validation and testing, the entire model is used, leading to a higher accuracy/lower loss.
@@
Figure 7(a) shows the test accuracy of the model versus SNR for simulated data. Figure 7(b) shows the test accuracy of the model versus SNR for hardware-captured data. In both cases, the test dataset for each SNR contains a combination of all possible values of $N_{UE}=\{0,1,2,\dots,12\}$. Both Figures 7(a) and 7(b) show that model accuracy increases with SNR. For the simulated data, the model significantly outperforms the DFT algorithm for all values of $\Delta$.
@@
For the hardware captured data, when $\Delta=0$, the model performs slightly better than the DFT algorithm. But as the SNR increases, the accuracies of the model and that of the DFT algorithm converge. This is because $\Delta=0$ constitutes a best-case scenario in which both algorithms are aided by the fact that the receiver knows exactly how many UEs are multiplexed. At higher values of $\Delta$, the convergence is not seen, and the NN model shows a significant gain compared to the DFT algorithm.
@@
Furthermore, from Figures 7(a) and 7(b), we observe an interesting difference between the performance of the neural network with simulated data versus hardware captured data. For all values of $\Delta$, there is a significant SNR gain with the hardware data compared to simulated data. For example, when $\Delta=0$, the accuracy with hardware captured data reaches 100% at as low an SNR as $5$ dB. In contrast, the accuracy with simulated data requires more than 1515dB SNR to reach 100%. This difference could be attributed to the characteristics of the datasets themselves. The simulated dataset includes a strong fading component that impairs the transmitted signal. However, owing to the short propagation distance, such a strong fading component is absent in the current 5G testbed setup that was used to obtain the hardware captures. We note that the captures still incorporate other real-world impairments caused by front-end hardware elements like filters, DACs, ADCs, up/down converters, and amplifiers. In summary, the hardware dataset models impairments due to front-end elements well. It also models over-the-air impairments but does not encapsulate all characteristics of a typical communication link between a gNB and UE that might span hundreds of meters. On the other hand, the simulated data models over-the-air characteristics well but it doesn’t model impairments due to hardware elements. We also point out that the IIT Madras 5G Testbed is moving towards deployment of Remote Radio Heads at on-campus sites which will allow the inclusion of even more realistic field data into future work.
@@
Another observation from Figures 7(a) and 7(b) is that the DFT algorithm performs worse on hardware data, showing that it is negatively affected not only by impairments from hardware elements but also by even the slightest over-the-air fading. On the other hand, the NN Model outperforms the DFT method in all scenarios of hardware and simulated datasets.
@@
Figure 8(a) shows the test accuracy of the NN model versus the number of multiplexed UEs ($N_{UE}$) for simulated data. Figure 8(b) shows the test accuracy of the NN model versus the number of multiplexed UEs ($N_{UE}$) for hardware-captured data. In both cases, the test dataset for each $N_{UE}$ contains a combination of all possible values of SNR $\{0,5,10,15,20\}$ dB. It is clear that the accuracy decreases as the maximum offset $\Delta$ increases. Note that no value is plotted when $N_{UE}=0$ and $\Delta=0$ since this case corresponds to the scenario in which no UEs are transmitting on a given Resource Block and the receiver knows apriori that no UEs are scheduled as well. In this case, the receiver need not even run. When $N_{UE}=12$, all the multiplexed UEs transmit only Scheduling Requests (SR). For both the simulated and hardware datasets, the DFT algorithm selects all 12 values from the 12-point DFT and assigns one value to each UE. Since all UEs are transmitting SRs, the order of assignment doesn’t matter, and hence the accuracy is always 100%. The NN model also seems to recognize that the only scenario in which 12 UEs can be multiplexed is when they all transmit exactly the same information, i.e., SRs.
Lower values of $N_{UE}$ tend to result in lower accuracy values. This is because the DFT algorithm picks up the top $N_{UE}$ values of the 12-point DFT, and the probability of a true value lying in one of the $12-N_{UE}$ locations increases as $N_{UE}$ decreases. We observe a similar pattern in the case of decoding with the proposed UCINet0 as well.
@@
Figure 7: Testing Accuracy vs SNR for (a) simulated data and (b) hardware captured data. The test dataset for each SNR contains a combination of $N_{UE}=\{0,1,2,\dots,12\}$.
@@
Figure 8: Testing Accuracy vs Number of Multiplexed UEs ($N_{UE}$) for (a) simulated data and (b) hardware captured data. The test dataset for each value of $N_{UE}$ contains a combination of SNR = $\{0,5,10,15,20\}$ dB.
IV-C Confusion Matrices and Column Charts

### v2

@@ V-B REAL-TIME OVER-THE-AIR DATASETS (LAB SETUP)
In addition to simulation datasets, we test the model using real-time over-the-air captures to assess generalizability. We use hardware captures derived from the state-of-the-art 5G testbed at IIT Madras [6] for more realistic testing. One of the hardware setups used (see Figure 7) consists of an N5182B Vector Signal Generator (VSG) for transmitting the 5G signal at a center frequency of $3.49986$ GHz (one of the sub-6 GHz channel raster in the n78 band). In our setup, the VSG acts like a UE, and it uses a commercial omnidirectional wideband monopole antenna to transmit the signals. The VSG is connected to the antenna through 2 SMA cables with $1.9$ dB wire loss each.
@@
A multi-channel 5G Remote Radio Head (RRH) with a dual-polarized antenna receives the signals over the air (for the purpose of the paper, we utilize only one antenna and one transceiver chain). The RRH operates in the n78 band with $100$ MHz bandwidth. It is ORAN-compliant and follows the 7.2b split as defined in [16]. Other receiver components of the RRH include an in-house Low Noise Amplifier (LNA) with 60dB gain at the receiver front end and an ADRV 9009 RF transceiver. In this indoor lab setup, we place the transmitter antenna one meter away from the receiver. The PUCCH Format 0 signal is transmitted from the VSG through its antenna, over the air, and then received at the RRH antenna, followed by the LNA and the transceiver. The signal from the 16-bit ADC is then collected and used for testing.
@@
Figure 7: IIT Madras 5G testbed lab setup with the Remote Radio Head used as a receiver (gNB) and the VSG used as a transmitter (Emulated UE).
V-C REAL-TIME OVER-THE-AIR DATASETS (LIVE NETWORK)
In addition to the indoor lab setup, we also incorporate data captures obtained from an outdoor RRH receiving PUCCH Format 0 signals from a commercial UE (Samsung A23 with a Qualcomm Snapdragon 695 modem). This RRH is part of the 5G network being run at IIT Madras in the n78 frequency band with custom developed hardware and covers the academic area of the campus.
The goal of this setup is to capture the PUCCH Format 0 signaling between a UE and gNB as part of a real 5G NR link, as opposed to a pre-loaded signal transmitted by the VSG in the lab setup, above. The RRH is part of an indigenous 5G NR gNB deployed on the IIT Madras campus, and is compliant to 3GPP and ORAN 7.2 split-standards. The RRH is the front-end responsible for the transmission and reception of signals. It also performs some signal processing functions like precoding, Fast Fourier Transforms (FFTs) and filtering. The other component of the gNB, the Baseband Unit (BBU), is responsible for functions like Polar Coding, Low Density Parity Check (LDPC) Coding, Channel Estimation, Channel Equalization, Cyclic Redundancy Check (CRC). Decoding of the information bits present in 5GNR control and data channels including PUCCH Format 0, is also done at the BBU. Note that, in out setup, these physical layer algorithms of the BBU are implemented on a PCIe form factor FPGA card.
Figure 8: IIT Madras Live 5G campus Network: (a) Outdoor RRH (16/32 channel MIMO systems) and (b) Commercial Samsung UE connected to a laptop running QXDM.
To facilitate PUCCH Format 0 data capture, first, we enable data logging at both the UE and the gNB. Second, we initiate the latching procedure by which the UE decodes the Synchronization Signal Burst (SSB) and System Information Block 1 (SIB1) transmitted by the gNB. The UE then identifies itself to the gNB by transmitting a PRACH signal. This triggers the rest of the initial access procedure including both uplink and downlink control and data channel signaling, including PUCCH Format 0.
V-C1 UE SIDE DATA (GROUND TRUTH) COLLECTION
Qualcomm eXtensible Diagnostic Monitor (QXDM) installed on a laptop connected to the UE enables collection of L1, L2 and L3 signaling logs. The saved logs are filtered to obtain only the NR5G MAC UCI Payload Information which contains PUCCH Format 0 specific parameters such as Num HARQ bits, Num SR bits, HARQ Payload and SR Payload. These parameters can be used to generate the ground truth labels for each Format 0 instance. The logs also capture parameters such as System Frame Number (SFN), Slot Number, Start Symbol, Num Symbol which can be used to synchronize the UCI payloads with the Format 0 signals captured at the gNB receiver.
V-C2 gNB SIDE DATA COLLECTION
We collect the frequency domain data received at the BBU along with the corresponding configuration parameters including SFN, Slot Number, number of scheduled UEs ($N_{UE}$) and Num HARQ bits, Num SR bits for each UE. The DFT-based correlation algorithm implemented on the BBU then decodes the UCI content from the frequency domain data. All of this information is packetized and sent to an off-the-shelf server via the PCIe interface at regular intervals. At the server, we write the data received from this PCIe interface to memory for further processing.
V-C3 DATA SYNCHRONIZATION
It is important to synchronize the data from two independent sources - QXDM and BBU logging. This synchronization happens in two steps: 1) Coarse synchronization with the system time stamp at both ends, 2) Fine synchronization is achieved based on the SFN and Slot Number embedded in the packets. Once synchronized, we can map the ground truth labels (UCI transmitted by the UE) to the appropriate PUCCH Format 0 signal received by the gNB and the UCI content decoded by DFT-based conventional algorithm (for comparison).
VI MODEL TRAINING AND TESTING
In this section, we describe the model training and testing process including forward and backward propagation, optimizer and loss, as well as the specific datasets used for training and testing. The training and testing were done using Python and TensorFlow on an NVIDIA A100 Tensor Core GPU. The code is available at [15] for download.
VI-A MODEL TRAINING
During training, we employ dropout [17] with a probability of 0.5, for all the models described in this paper. Dropout achieves both ensemble learning [18] and regularization [19] by “dropping out” or cutting off certain neurons and their connections randomly with a certain probability. During each forward pass, the sigmoid activation at the output layer generates probabilities which are used to compute a binary cross entropy loss. The loss is then back-propagated using Stochastic Gradient Descent (SGD) with a learning rate of $10^{-2}$ and a momentum parameter of $0.9$.
Based on the insights gained from our previous work [5], we train the neural network using data with a middle SNR of $10$ dB. We have found that this model achieves good generalization across the entire range of SNR during testing, i.e., $10$ dB SNR is sufficient enough for the model to learn the true patterns both in the underlying data as well as noise. Training data also includes a $\Delta$ value of $2$ (middle value of metadata offset), all UE multiplexing scenarios ($\tilde{N}_{UE}=0,1,\dots 12$), the TDLC300 channel model, and Doppler frequency shifts of $0,500,1000,1500$, and $2000$ Hz. $75\%$ of the data is used for training, from which a further 30% is used for validation and fine-tuning of hyperparameters. We train the model for 150 epochs (an epoch is 1 pass of the entire dataset through the neural network). The training dataset size is $1638K$.
VI-B MODEL TESTING
To ensure consistent performance of the model, we test it across a wide range of parameters using both simulated and hardware datasets. For the simulated data, the testing instances span SNRs of $0,2,4,\dots 20$dB, TDLC300, TDLC30, TDLA300, TDLA30 channel models, Doppler frequency shifts of $0,400,800,1200,1600$ and $2000$ Hz, all UE multiplexing scenarios ($\tilde{N}_{UE}=0,1,\dots 12$) and $\Delta$ values of 0, 22 and $4$. For the hardware data, the testing instances span SNRs $0,5,10,15,$ and $20$ dB, all UE multiplexing scenarios ($\tilde{N}_{UE}=0,1,\dots 12$) and and $\Delta$ values of 0, 22 and $4$. For the live network data, a single UE connected to the QXDM tool is subjected to different SNRs ($1$, 33, 77, and $10$ dB), and the data is collected at the gNB receiver.
Based on the known scheduling information from L2, the superset of all possible cyclic shifts a UE can transmit can be pre-determined, as shown by the $\overline{\alpha}$ examples in Section II-C. Even in the case of a multiplexed UE scenario, the possible outputs of the neural network (alpha indicators) can be derived by combining all the individual UE’s possible cyclic shifts. This information, in the form of a binary mask vector is applied to the outputs of the models, during inference. The function of the mask is to nullify the specific cyclic shift predictions that are apriori known to be incorrect. We also observe a significant gain when the inference mask is applied to the neural network’s prediction.
For example, when $2$ UEs with each of them being scheduled with a $1$ SR and $1$ HARQ transmission, a total of $4$ cyclic shifts could be transmitted individually. So, the binary mask will assert all the $8$ possible cyclic shifts as the inference mask to the output of the neural network.
VII RESULTS AND DISCUSSION
@@
In this section, we describe the performance results of the UCINet0 NN architecture shown in Figure 3.
VII-A PERFORMANCE METRICS AND THEIR MOTIVATION
@@
We use model test accuracy as the main performance metric. We also show confusion matrices to indicate the distribution of model predictions. Wherever appropriate, we show comparisons of the model performance with the corresponding performance shown by the DFT-based correlation algorithm.
@@
Accuracy is simply the number of correct predictions in relation to total predictions. Though it is a useful high-level metric to gauge model performance, in cases where the classes in a dataset are unequally distributed or when, as is the case in this paper, there are more than 2 classes, accuracy alone as a metric for an NN classifier does not offer the full picture of model performance. It is helpful to know if the model classifies all classes equally well or if it is more “confused” by certain classes compared to others. Hence, calculating a confusion matrix gives a better insight into any patterns that may exist in both correct and wrong classifications made by the model. In this paper, we show 3 types of confusion matrices: (1) A multi-label confusion matrix that shows how well each $\alpha$ is predicted; (2) A confusion matrix showing how well the model predicts the correct number of multiplexed UEs, and (3) A confusion matrix showing how each value of $\alpha$ (including the cases when no $\alpha$ is selected) is classified. We present this last metric in the form of a column chart for ease of visualization. These 3 types of confusion matrices provide a sense of interpretability of the inner workings of the NN model.
VII-B ACCURACY AND LOSS
@@
Figure 9(a) shows the training and validation accuracy with respect to the training epochs. Figure 9(b) shows the training and validation loss. Both curves follow the expected pattern for a well-fit model. The accuracy gradually increases and the loss steadily decreases with epoch and eventually converges. We also note that the validation accuracy and loss values are slightly better than the corresponding training values. One possible explanation for this could be the use of dropout. As stated above, a dropout with probability 0.5 drops neurons $50\%$ of the time during training. However, during validation and testing, the entire model is used, leading to a higher accuracy/lower loss.
@@
Figure 10(a) shows the test accuracy of the model versus SNR for simulated data. Figure 10(b) shows the test accuracy of the model versus SNR for hardware-captured data. In both cases, the test dataset for each SNR contains a combination of all possible values of $\tilde{N}_{UE}=\{0,1,2,\dots,12\}$. Both Figures 10(a) and 10(b) show that model accuracy increases with SNR. For the simulated data, the model significantly outperforms the DFT-based correlation algorithm for all values of $\Delta$.
@@
For the hardware-captured data, when $\Delta=0$, the model performs slightly better than the DFT-based correlation algorithm. However, as the SNR increases, the accuracy of the model and that of the DFT-based correlation algorithm converge. This is because $\Delta=0$ constitutes a best-case scenario in which both algorithms are aided by the fact that the receiver knows exactly how many UEs are multiplexed. At higher values of $\Delta$, the convergence is not seen, and the NN model shows a significant gain compared to the DFT-based correlation algorithm.
@@
Furthermore, from Figures 10(a) and 10(b), we observe an interesting difference between the performance of the neural network with simulated data versus hardware-captured data. For all values of $\Delta$, there is a significant SNR gain with the hardware-captured data compared to simulated data. For example, when $\Delta=0$, the accuracy with hardware-captured data reaches 100% at as low an SNR as $5$ dB. In contrast, the accuracy with simulated data requires more than 1515dB SNR to reach 98%). This difference could be attributed to the characteristics of the datasets themselves. The simulated dataset includes a strong fading component (both slow and fast) that impairs the transmitted signal. However, owing to the short propagation distance and stationary VSG, such a strong fading component is absent in the current 5G testbed setup that was used to obtain the hardware captures. We note that the captures still incorporate other real-world impairments caused by front-end hardware elements like filters, DACs, ADCs, up/down converters, and amplifiers. In summary, the hardware dataset models impairments due to front-end elements well. It also models over-the-air impairments but does not encapsulate all characteristics of a typical communication link between a gNB and UE that might span hundreds of meters. On the other hand, the simulated data models over-the-air characteristics well, but it doesn’t model impairments due to hardware elements. We also point out that the IIT Madras 5G Testbed is moving towards a full-fledged deployment of Remote Radio Heads at on-campus sites, which will allow the inclusion of even more realistic field data into future work.
@@
Another observation from Figures 10(a) and 10(b) is that the DFT-based correlation algorithm performs worse on hardware data, showing that it is negatively affected not only by impairments from hardware elements but also by even the slightest over-the-air fading. On the other hand, the NN Model outperforms the DFT method in all scenarios of hardware and simulated datasets.
Figure 11 exhibits the model performance for the data captured from a live network. We observe that the model’s performance remains consistent across SNRs, reflecting the robustness of the model in live network signalling.
@@
Figure 12(a) shows the test accuracy of the NN model versus the number of multiplexed UEs ($\tilde{N}_{UE}$) for simulated data. Figure 12(b) shows the test accuracy of the NN model versus the number of multiplexed UEs ($\tilde{N}_{UE}$) for hardware-captured data. In both cases, the test dataset for each $\tilde{N}_{UE}$ contains a combination of all possible values of SNR $\{0,5,10,15,20\}$ dB. It is clear that the accuracy decreases as the maximum offset $\Delta$ increases. Note that no value is plotted when $\tilde{N}_{UE}=0$ and $\Delta=0$ since this case corresponds to the scenario in which no UEs are transmitting on a given Resource Block and the receiver knows apriori that no UEs are scheduled as well. In this case, the receiver need not even run. When $\tilde{N}_{UE}=12$, all the multiplexed UEs transmit only SRs. For both the simulated and hardware datasets, the DFT-based correlation algorithm selects all 12 values from the 12-point DFT and assigns one value to each UE. Since all UEs are transmitting SRs, the order of assignment doesn’t matter, and hence the accuracy is always 100%. The NN model also seems to recognize that the only scenario in which 12 UEs can be multiplexed is when they all transmit exactly the same information, i.e., SRs.
@@
Figure 10: Testing Accuracy vs SNR for (a) simulated data and (b) hardware-captured data (lab setup). The test dataset for each SNR contains a combination of $\tilde{N}_{UE}=\{0,1,2,\dots,12\}$
Figure 11: Testing Accuracy vs average SNR (approximate) for data captured from the live network. These average SNRs have been obtained at different locations of the network.
@@
Figure 12: Testing Accuracy vs Number of Multiplexed UEs ($\tilde{N}_{UE}$) for (a) simulated data and (b) hardware-captured data. The test dataset for each value of $\tilde{N}_{UE}$ contains a combination of SNR = $\{0,2,4,\dots,20\}$ dB for (a) and SNR = $\{0,5,10,15,20\}$ dB for (b).
Figure 13 shows the test accuracy of the NN model versus SNR, for simulated data, across multiple Doppler frequency shifts. The graph compares both the NN and the conventional DFT-based correlation algorithm. We observe that across SNRs, for any Doppler frequency shift, the neural network based decoder performs consistently better than the conventional approach. The variation in accuracy values for different Doppler frequency shifts is minimal for the NN based decoder when compared to the conventional decoder.
Figure 13: Accuracy versus SNR graph for multiple Doppler frequency shifts ranging from 0 to 2000 Hz.
Figure 14 shows the test accuracy of the NN model versus SNR, for simulated data, across multiple channel models. The graph compares both the NN and the conventional DFT-based correlation algorithm. Noting that the model has been trained exclusively on the TDLC300 channel model, we observe that for delay spreads lower than 300 ns, for both TDLA and TDLC, the performance is equivalent to that of TDLC300. For TDLA300, there is a performance drop of approximately $1\%$ across all SNRs. While this can be considered to be an acceptable generalization performance across various channel models, the $1\%$ gap can easily be reduced by including some instances of TDLA300 in the training.
Figure 14: Accuracy vs SNR for multiple fading channel scenarios
VII-C CONFUSION MATRICES AND COLUMN CHARTS
In this subsection, confusion matrices and column charts are presented to gain deeper understanding of the model’s performance. Considering the relatively low variance in accuracy across various Doppler frequency shifts, as shown in Figure 13, further analysis incorporates all the Doppler frequency shifts together in the same dataset.